OpenStudy (anonymous):

The mean of a set of data is fifteen. If there are five values in the data set, what is the sum of the values ?

OpenStudy (anonymous):

@saifoo.khan

OpenStudy (saifoo.khan):

\[\frac{a+b+c+d+e}{5} = 15\]

OpenStudy (anonymous):

mhmm. how do you find the values?

OpenStudy (anonymous):

oh, right.

OpenStudy (anonymous):

one moment

OpenStudy (anonymous):

they have to equal 75?

OpenStudy (anonymous):

a-e are all 75,yes?

OpenStudy (saifoo.khan):

Yes sir.

OpenStudy (anonymous):

Call me '123'.

OpenStudy (anonymous):

Ok, so 75+75+75+75+75/5=15?

OpenStudy (saifoo.khan):

I picked up my cell, dialed 123 but it says the number is incomplete. :P

OpenStudy (anonymous):

Lol!

OpenStudy (saifoo.khan):

NO! it can't be 75+75+75+75+75/5=15?

OpenStudy (anonymous):

Okay,okay.

OpenStudy (anonymous):

I'm sorry. Let me try again.

OpenStudy (saifoo.khan):

Sure.

OpenStudy (anonymous):

And I meant my nickname is one hundred and twenty three. Or 123 for short:)

OpenStudy (saifoo.khan):

Ohh. Sir 123.

OpenStudy (anonymous):

Oh -- is it \[3+3+3+3+3/5=15?\]

OpenStudy (anonymous):

Indeed. But that would Have to be 'Ms.'

OpenStudy (anonymous):

Never mind, sorry, my reply above is wrong.

OpenStudy (saifoo.khan):

NOEE!!!! you can't assume any numbers. It will jsut be a+b+c+d+e.

OpenStudy (anonymous):

Oh! So I just write it as that? a+b+c+d+e/5=15?

OpenStudy (saifoo.khan):

a+b+c+d+e = 75

OpenStudy (saifoo.khan):

Simplify it.

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

I really dont understand this one....srry!

OpenStudy (anonymous):

It;'s ok.

OpenStudy (saifoo.khan):

@darkb0nebeauty : Which part you don't get?

OpenStudy (anonymous):

just knw i cant help her..so sorry! @saifoo.khan

OpenStudy (anonymous):

@timo86m

OpenStudy (saifoo.khan):

Just simply multiply 5 on both sides.

OpenStudy (anonymous):

Hmm.....

OpenStudy (anonymous):

Then y want u do it if u now it! @saifoo.khan

OpenStudy (anonymous):

the sum is 75

OpenStudy (anonymous):

Ah, okay, thanks,sir!

OpenStudy (anonymous):

sum of a set/5=15 solve for sum of the set. sum of the set is a+b+c+d+e it doesn't ask for individual variables but the whole sum of the set.